OpenStudy (anonymous):

"suppose \(c|(a+b)\) where \(a,b,c\in\mathbb{Z}\). then \(c|(pa+qb)\) where \(p,q\in\mathbb{Z}\)." can you really make such assumption!?

OpenStudy (anonymous):

"for SOME integers \(p\) and \(q\)" i meant to add

OpenStudy (anonymous):

i don't seem to get why x.x

OpenStudy (jamesj):

Choose p=q=1

OpenStudy (anonymous):

No we can't make such assumption.

OpenStudy (anonymous):

like a=2, b=4, c=3 c|(a+b) -> 3|6 which is true but c|(2a+3b) -> 3|16 is false :(

OpenStudy (jamesj):

\[ c | (a+b) \implies c | (1\cdot a + 1\cdot b) \]

OpenStudy (anonymous):

\(3|(4+5)\) but 3 does not divide \((4\times 3+5 \times 5) \)

OpenStudy (anonymous):

ooooh

OpenStudy (jamesj):

so if the question is there exist at least one pair p,q such your statement is true, then yes.

OpenStudy (anonymous):

But this question is asking for for all \(p,q \in \mathbb{Z}\), hence incorrect.

OpenStudy (anonymous):

all or any*

OpenStudy (jamesj):

ffm, pre-algebra says immediately below "for SOME p and q ..."

OpenStudy (anonymous):

Hey, EDIT feature is a must!!!!

OpenStudy (anonymous):

I have a tendency not to read any comment/answer before trying it on my own.

OpenStudy (jamesj):

no kidding

OpenStudy (anonymous):

"for SOME integers p and q" that's true.